OpenStudy (anonymous):

Solve x^2 – 6x = 27 by completing the square. Solve x^2 – 6x = 27 by completing the square. @Mathematics

OpenStudy (anonymous):

(x-3)^2 - 9 = 27-9

OpenStudy (hoblos):

(x-3)^2 = 36 x-3 = +/- 6 => x=9 or x = -3

OpenStudy (anonymous):

oopppsss

OpenStudy (anonymous):

i put too many -9 ' s